A mixture of 10.325 g of iron (III) oxide and 5.734 g of aluminum metal is placed in a crucible and heated in a high-temperature oven, where iron and aluminum oxide is formed. How many atoms of iron is formed during this reaction? (2 pts) How much aluminum, in grams, is left over after the chemical reaction has occurred? (2 pt) a.
Answer #1

Fe2O3 + 2Al -----------> 2Fe + Al2O3

according to balanced reaction

159.69 g Fe2O3 gives 2 x 6.023 x 1023 atoms of Fe

10.325 g Fe2O3 gives 10.325 x 2 x 6.023 x 1023 / 159.69 = 7.80 x 1022 Fe atoms

atoms of Fe formed = 7.80 x 1022 atoms

a) 159.69 g Fe2O3 reacts with 2 x 27 g Al

10.325 g Fe2O3 reacts with 10.325 x 2 x 27 / 159.69 = 3.491 g Al

mass of Al left unreacted = 5.734 - 3.491 = 2.243 g

mass of Al left = 2.243 g